OpenStudy (anonymous):

Find all the real-number roots of the equation. Give an exact expression for the root and also a calculator approximation rounded to three decimal places. log3 x + log3(x + 8) = 0

OpenStudy (anonymous):

\[\log_3(x)+\log_3(x+8)=0\] \[\log_3(x(x+8))=0\] \[3^0=x(x+8)\] \[x(x+8)=1\]

myininaya (myininaya):

\[\log_3(x(x+8))=0\] 3^ 3^ \[x(x+8)=1\] \[x^2+8x-1=0\] \[x=\frac{-8 \pm \sqrt{64-4(1)(-1)}}{2}=\frac{-8 \pm \sqrt{68}}{2}\]

OpenStudy (anonymous):

\[x^2+8x-1=0\] \[x=-4+\sqrt{17}\]

OpenStudy (anonymous):

what is this strange formula i see above?

myininaya (myininaya):

but there is only one x

OpenStudy (anonymous):

yes indeed there is only one x, right between w and y

myininaya (myininaya):

since one is will make the inside of the log above negative

OpenStudy (anonymous):

\[\color{pink}{\text{you are getting sleepy... very sleepy}}\]

myininaya (myininaya):

\[x=\frac{-8 + \sqrt{68}}{2}\] i trust you can simplify

OpenStudy (anonymous):

@tky myininaya got the same answer. there are two roots of \[x^2+8x-1=0\] but one is negative and since the domain log is positive numbers you cannot use the negative solution

OpenStudy (anonymous):

and yes, i can simplify \[\frac{-8+\sqrt{68}}{2}\] by never getting it to begin with \[x^2+8x-1=0\] \[x^2=8x-1\] \[(x+4)^2-1+16=17\] \[x+4=\pm\sqrt{17}\] \[x=-4\pm\sqrt{17}\]

OpenStudy (anonymous):

typos galore, which is what i get for being a wise guy. should be \[x^2+8x=1\] \[(x+4)^2=1+16=17\] etc